solve the following system of equations using substiution. Please help!
x-4y=4
6x-3y=-18

Answer by mananth(16946)

1 x -4 y = 4 .............1
-3 y = -18 -6 x ---------2
/ -3 =
y = 6 + 2 x
Substitute y in (1)
1 x -4 ( 6 + 2 x ) = 4
1 x -24 + -8 x = 4
1 x -8 x = 4 24
-7 x = 28
/ -7
x= -4
Plug the value of x in (1)
1 x -4 y = 4
1 * -4 -4 y = 4
-4 -4 y = 4
-4 y = 8
/ -4
y= -2
m.ananth@hotmail.ca